# J-PLUS: Spectroscopic validation of H emission line maps in spatially resolved galaxies

### v1

@@ Abstract
Aims. We present a dedicated automated pipeline to construct spatially resolved emission H+[NII] maps and to derive the spectral energy distributions (SEDs) in 12 optical filters (five broad and seven narrow/medium) of H emission line regions in nearby galaxies (z 0.0165) observed by the Javalambre Photometric Local Universe Survey (J-PLUS).
Methods. We used the filter of Å width centered at Å to trace H + [NII] emission and and broad bands were used to estimate the stellar continuum. We create pure emission line images after the continnum subtraction, where the H emission line regions were detected. This method was also applied to Integral Field Unit (IFU) spectroscopic data from PHANGS-MUSE, CALIFA and MaNGA surveys by building synthetic narrow-bands based on J-PLUS filters. The studied sample includes the cross-matched catalog of these IFU surveys with J-PLUS third data release (DR3), amounting to PHANGS-MUSE, CALIFA, and MaNGA galaxies at , respectively.
@@
Conclusions. The J-PLUS data provides reliable spatially resolved H+[NII] emission maps for nearby galaxies. We provide the J-PLUS DR3 catalog for the galaxies with IFU data, including emission maps, SEDs of star-forming clumps, and radial profiles.
@@
Spatially resolved studies of galaxies have significantly advanced our understanding of galaxy formation and evolution. Unlike previous techniques that gather integrated flux from entire galaxies, spatially resolved studies reveal local variations in different regions of galaxies, allowing the detailed understanding of different characteristics such as stellar population properties, star formation, gas kinematics, and the chemical composition in galaxies. In particular, spatially resolving galaxies has provided new opportunities for detailed emission line studies in galaxies. For example, H (6563 Å) is a prominent emission line in the optical, used as an important star formation tracer ( 10 Myr, Kennicutt, 1998) , which is produced from hydrogen recombination in nebular regions ionized by the radiation field of young, short-lived massive stars, and also due to the contribution from (AGN). Spatially resolved H emission line maps (hereafter ”emap”) are crucial for understanding the star formation activity, gas dynamics, and feedback processes in galaxies.
@@
Narrow-band (NB) imaging and integral field spectroscopy (IFS) are two key techniques that have allowed spatially resolved emission line studies of galaxies. The NB technique dedicated for H imaging surveys (e.g., James et al., 2004; Sobral et al., 2013; Stroe & Sobral, 2015; Sobral et al., 2015) has greatly improved our understanding of spatially resolved information on star formation in galaxies. However, these post NB surveys were usually limited to a few square degrees or a selected number of galaxies (e.g., 100 to 1000).
@@
The introduction of IFS in astronomy has been a game changer for spatially resolved studies of galaxies, enabling astronomers to simultaneously gather spectra from different regions of a galaxy in the form of 3D data cubes. The TIGER instrument (Bacon et al., 1988) developed at the Observatoire de Lyon, France mounted on the Canada-France-Hawaii Telescope (CFHT) was the first instrument that demonstrated the power of IFS in astronomy. Then, more advanced optical Integral Field Unit (IFUs) like SAURON (Miller et al., 2000), VIMOS (Le Fèvre et al., 2003), SAMI (Croom et al., 2012), MUSE (Bacon et al., 2014), MaNGA (Bundy et al., 2015), and PPaK/CALIFA (Sánchez et al., 2012) were introduced into the field and these have significantly improving our understanding of galaxies, nebula, star-forming regions by offering a resolved perspective that allows for detailed analysis of their internal structures and physical processes. The earlier IFS systems such as SAURON or Tiger were lenslet- based, MaNGA and PPaK/CALIFA are fiber-based IFS, while MUSE is image slicer-based. Recently, IFUs have expanded to space-based missions such as the Near-Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022) on the James Webb Space Telescope (JWST), enabling unprecedented high-resolution spectroscopy of distant galaxies. These IFSs have studied H emission line maps of thousands of galaxies that allow for statistical studies of galaxy evolution, environmental effects, and the relationship between star formation and galaxy morphology (Novais & Sodré, 2019; Barrera-Ballesteros et al., 2023). However, these IFUs are time-expensive and limited due to the small field of view (FoV) and limited number of observed targets.
@@
Multiband photometric surveys such as ALHAMBRA (Advanced Large, Homogeneous Area Medium-Band Redshift Astronomical Survey; Moles et al., 2008) covering 4 deg2, PAU (Physics of the Accelerating Universe; Benítez et al., 2009) covering 100 deg2 have given a significant contribution to wide-field surveys and bridge the gap between broad-band photometry and IFS with their multiple filters. Javalambre Photometric Local Universe Survey (J-PLUS; Cenarro et al., 2014, 2019) and Javalambre-Physics of the Accelerating universe astrophysical Survey (J-PAS; Benitez et al., 2014; Dupke et al., 2016) are the two ongoing optical surveys in the northern sky at the Observatorio Astrofísico de Javalambre (OAJ). The large FoV, high spatial resolution and contiguous narrow-band filters covering a wide wavelength range(3300-11000 Å) of J-PAS act like a low-resolution IFU (R 60), which is suitable for IFU-like science such as spatially resolved studies of galaxies planned over 8500 deg2 (by the end of 2030). J-PLUS includes some of the narrow bands from J-PAS covering prominent emission lines such as [OII], H+[NII] at z¡0.0165 and also [OIII] at 0.006¡z¡0.0473 (Lumbreras-Calle et al., 2022) allowing us to study the spatially resolved properties of galaxies in these spectral lines. In comparison to the field of view of other time-expensive IFUs such as MaNGA (12′′- 32′′), SAMI (15′′), PPaK/CALIFA (60′′), and MUSE (60′′), the uniqueness of J-PLUS is their large FoV () offering a large contiguous observing area to understand the resolved structure of galaxies, galaxy groups, galaxy clusters and thus, probe their environment.
@@
The method of measuring H fluxes and SFRs from J-PLUS photometry was first introduced by Vilella-Rojo et al. (2015) and Logroño-García et al. (2019). Vilella-Rojo et al. (2021) applied their method to estimate the star formation main sequence in the local Universe using J-PLUS DR1. Here, we aim to extend that work to validate the H flux extraction for spatially resolved studies using J-PLUS DR3. We selected galaxies with redshifts up to z 0.0165, ensuring that the J0660 NB filter covers the H emission line within this range. This leads to thousands of galaxies at D 70 Mpc within the 3192 deg2 observed area of J-PLUS DR3. At this redshift range, J-PLUS will be able to resolve star-forming regions on the physical scale of several hundred parsecs ( 168 pc). We validate our method of building the emaps by comparing them with other well-known IFU observations of the same galaxies. This is the first of a series of works where we aim to use the H map to study the spatially resolved star formation main sequence of J-PLUS galaxies and UV and H studies of J-PLUS galaxies.
@@ 2.1 J-PLUS DR3
J-PLUS (Cenarro et al., 2014, 2019) is a multi-band photometric survey designed to observe the northern sky in several thousand square degrees using the 80cm Javalambre Auxiliary Survey Telescope (JAST80) at OAJ, Teruel. It has 12 broad, intermediate, and narrow band optical filters with 2 deg2 FoV. The properties of J-PLUS filters are given in Table 1. After two data releases (DR1 on July 2018, and DR2 on July 2020), J-PLUS had its third data release in July 2022, containing 1642 J-PLUS fields covering 3192 deg2. The pixel scale of J-PLUS is 0.555′′/pixel. The point spread function (PSF) of J-PLUS observation is 1.1 arcsecs in the r band. The DR3 images that we have used here are available from the OAJ database 111https://archive.cefca.es/catalogues/jplus-dr3. The calibration and zeropoints are taken from López-Sanjuan et al. (2019, 2024).
@@ 2.2 PHANGS-MUSE
The Multi-Unit Spectroscopic Explorer (MUSE; Bacon et al., 2010) is an IFS at the Very Large Telescope (VLT) of the European Southern Observatory (ESO) Paranal Observatory, Chile. It is composed of 24 identical integral-field units (IFUs). In its wide field mode (WFM; FoV of ) 90000 spectra are taken with a spatial resolution of 0.2′′/pixel covering a wavelength range of 4750 - 9350 Å with a spectral resolution of 2.5 Å. The spectral resolution of MUSE is 2.5 Å(FWHM) at 7000 Å and the spectral sampling of 1.25Å /slice.
@@ 2.3 CALIFA
The Calar Alto Legacy Integral Field Area Survey (CALIFA; Sánchez et al., 2012) is a large project to obtain spatially resolved spectra of 600 nearby galaxies with a size of 64 ′′ (diameter) in the redshift range of 0.005 z 0.03 using a PMAS/PPAK integral field spectrophotometer, mounted on the Calar Alto 3.5 m telescope. It utilises an hexagonal field-of-view (FoV) of . It has two observing modes: Intermediate spectral resolution (V1200, R1550) and low resolution (V500, R850). The CALIFA observations were carried out between 2010 and 2016.
We used version 2.3 from the extended data release (eDR) of CALIFA (Sánchez et al., 2023) that includes 895 galaxies. This version has significantly improved spatial resolution and image quality. The PSF is improved from 2.4 arcsec to 1 arcsec of FWHM. The configuration includes the low-resolution setup (V500, R850), integration time of 900 s per pointing, with a wavelength range of 3745 and 7500 Å . The spatial sampling of the eDR CALIFA data is 0.5′′/pixels with a wavelength resolution of 2Å /slice. The eCALIFA cubes are downloaded from the CALIFA website 333https://ifs.astroscu.unam.mx. The downloaded CALIFA cubes are already corrected for Galactic extinction using the method by Schlegel et al. (1998) and Cardelli et al. (1989).
@@ 2.4 MaNGA
Mapping Nearby Galaxies at Apache Point Observatory (MaNGA; Bundy et al., 2015) is a program under the fourth phase of the Sloan Digital Sky Surveys (SDSS-IV; Blanton et al., 2017). MaNGA has mapped the kinematics and composition of 10,100 nearby galaxies in the redshift range of 0.01 z 0.15 with an average of z0.03 using 17 hexagonal shaped IFUs mounted on Sloan Foundation 2.5m telescope at Apache Point Observatory from 2014 to 2020. The IFU size ranges from 12 arcsecs (19 fibers) diameter to 32 arcsecs (127 fibers) with a wavelength coverage of 3600 - 10000 Å across 2700 deg2 and a spectral resolution of R2000. The medium spatial resolution of MaNGA is 2.54′′. The pixel scale of MaNGA is 0.5′′/pixel with a wavelength resolution of 0.834 Å /slice. The MaNGA cubes are downloaded from DR17 archive444https://data.sdss.org/sas/dr17/manga/spectro/redux/v3_1_1. The MaNGA fluxes are not corrected for Galactic extension.
@@ 3 Sample selection
There are 176 galaxies cross-matched with J-PLUS DR3 and eCALIFA, with 78 galaxies in the redshift range of z0.0165. Among them, 20 galaxies are early-type (E/SO), 9 galaxies show dust-lane features, and 48 galaxies show clear evidence of star formation activity (based on GALEX-UV and H + [NII] images) in the visual morphology and exhibit H emission line at 6562.8 Å rest frame wavelength in CALIFA spectra. Among the 48 galaxies, the [NII] line lies outside the filter window for redshifts greater than 0.014 for 10 galaxies. Above redshifts greater than 0.0165, the H line begins to fall outside the filter curve, and galaxies with such redshifts were excluded from the sample.
@@
The total sample includes 158 galaxies with 3D data cubes from either of the aforementioned IFUs. For the final analysis, we used 115 galaxies that exhibit H line in their spectra and have coverage of the full H line within the J0660 filter curve. Fig. 3 shows the coverage of J-PLUS DR3 with the distribution of the final sample selected from the cross-matched catalog. The information of the final cross-matched sample is given in Appendix LABEL:tab:catalog.
@@ 4 The J-SHE Pipeline for building emaps of J-PLUS and IFUs
In this section, we describe the methodology followed to build the emaps and radial profile of galaxies, SEDs, and spectra of emission line regions from J-PLUS and IFUs. The J-PLUS Spatially-resolved H Emission line (J-SHE) pipeline is written in GNU Make mainly utilizing GNU Astronomy Utilities (Gnuastro; Akhlaghi & Ichikawa, 2015) and Python packages. Gnuastro is an official GNU package for the manipulation and analysis of astronomical data. Our pipeline consists of three modules (see Fig. 4). The first and second modules are dedicated to building emaps for J-PLUS (hereafter, J-IFU module) and IFUs (hereafter, IFU module) respectively. The third module generates the galaxy radial profiles, SEDs, and spectra of emission line regions from the emaps (hereafter, Analysis module).

### v2

@@ Abstract
Aims. We present a dedicated automated pipeline to construct spatially resolved emission H+[NII] maps and to derive the spectral energy distributions (SEDs) in 12 optical filters (five broad and seven narrow and medium) of H emission line regions in nearby galaxies (z 0.0165) observed by the Javalambre Photometric Local Universe Survey (J-PLUS).
Methods. We used the filter of Å width centered at Å to trace H + [NII] emission, and and broad bands were used to estimate the stellar continuum. We created pure emission line images after the continnum subtraction, where the H emission line regions were detected. This method was also applied to integral field unit (IFU) spectroscopic data from PHANGS-MUSE, CALIFA, and MaNGA surveys by building synthetic narrow bands based on J-PLUS filters. The studied sample includes the cross-matched catalog of these IFU surveys with the J-PLUS third data release (DR3), amounting to two PHANGS-MUSE, CALIFA, and MaNGA galaxies at , respectively.
@@
Conclusions. The J-PLUS data provide reliable spatially resolved H+[NII] emission maps for nearby galaxies. We provide the J-PLUS DR3 catalog for the galaxies with IFU data, including emission maps, SEDs of star-forming clumps, and radial profiles.
@@
Spatially resolved studies of galaxies have significantly advanced our understanding of galaxy formation and evolution. Unlike previous techniques that gather an integrated flux from entire galaxies, spatially resolved studies reveal local variations in different regions of galaxies, allowing the detailed understanding of different characteristics such as stellar population properties, star formation, gas kinematics, and the chemical composition in galaxies. In particular, spatially resolving galaxies has provided new opportunities for detailed emission line studies in galaxies. For example, H (6563 Å) is a prominent emission line in the optical, used as an important star formation tracer ( 10 Myr, Kennicutt, 1998), which is produced from hydrogen recombination in nebular regions ionized by the radiation field of young, short-lived massive stars, and also due to the contribution from active galactic nuclei (AGN). Spatially resolved H emission line maps (hereafter ”emap”) are crucial for understanding the star formation activity, gas dynamics, and feedback processes in galaxies.
@@
Narrow band (NB) imaging and integral field spectroscopy (IFS) are two key techniques that have allowed spatially resolved emission line studies of galaxies. The NB technique dedicated to H imaging surveys (e.g., James et al., 2004; Sobral et al., 2013; Stroe & Sobral, 2015; Sobral et al., 2015) has greatly improved our understanding of spatially resolved information on star formation in galaxies. However, these post-NB surveys were usually limited to a few square degrees or a selected number of galaxies (e.g., 100 to 1000).
@@
The introduction of IFS in astronomy has been a game changer for spatially resolved studies of galaxies, enabling astronomers to simultaneously gather spectra from different regions of a galaxy in the form of 3D data cubes. The TIGER (Traitement Intégral des Galaxies par l’Étude de leurs Raies) instrument (Bacon et al., 1995) developed at the Observatoire de Lyon, France mounted on the Canada-France-Hawaii Telescope (CFHT) was the first instrument that demonstrated the power of IFS in astronomy. Then, more advanced optical integral field unit (IFUs) such as the Spectroscopic Areal Unit for Research on Optical Nebulae (SAURON; Miller et al., 2000), VIsible MultiObject Spectrograph (VIMOS; Le Fèvre et al., 2003), Sydney-AAO Multi-object Integral-field spectrograph (SAMI; Croom et al., 2012), The Multi Unit Spectroscopic Explorer (MUSE; Bacon et al., 2014), Mapping Nearby Galaxies at Apache Point Observatory (MaNGA; Bundy et al., 2015), and The PPaK Calar Alto Legacy Integral Field Area Survey (CALIFA; Sánchez et al., 2012) were introduced into the field and these have significantly improved our understanding of galaxies, nebula, and star-forming regions by offering a resolved perspective that allows for detailed analysis of their internal structures and physical processes. The earlier IFS systems such as SAURON or TIGER were lenslet-based, MaNGA and PPaK/CALIFA are fiber-based IFS, while MUSE is image slicer-based. Recently, IFUs have expanded to space-based missions such as the Near-Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022) on the James Webb Space Telescope (JWST), enabling unprecedented high-resolution spectroscopy of distant galaxies. These IFSs have studied H emission line maps of thousands of galaxies that allow for statistical studies of galaxy evolution, environmental effects, and the relationship between star formation and galaxy morphology (Novais & Sodré, 2019; Barrera-Ballesteros et al., 2023). However, these IFUs are time-expensive and limited due to the small field of view (FoV) and limited number of observed targets.
@@
Multiband photometric surveys such as the Advanced Large, Homogeneous Area Medium-Band Redshift Astronomical survey (ALHAMBRA; Moles et al., 2008) covering 4 deg2 and Physics of the Accelerating Universe (PAU; Benítez et al., 2009) covering 100 deg2 have significantly contributed to wide-field surveys and bridge the gap between broad-band photometry and IFS with their multiple filters. The Javalambre Photometric Local Universe Survey (J-PLUS; Cenarro et al., 2014, 2019) and Javalambre-Physics of the Accelerating universe astrophysical Survey (J-PAS; Benitez et al., 2014; Dupke et al., 2016) are the two ongoing optical surveys in the northern sky at the Observatorio Astrofísico de Javalambre (OAJ). The large FoV, high spatial resolution, and contiguous NB filters covering a wide wavelength range (3300-11000 Å) of J-PAS act as a low-resolution IFU (R 60), which is suitable for IFU-like science such as spatially resolved studies of galaxies planned over 8500 deg2 (by the end of 2030). J-PLUS includes some of the narrow bands from J-PAS covering prominent emission lines such as [OII] and H+[NII] at z¡0.0165 and also [OIII] at 0.006¡z¡0.0473 (Lumbreras-Calle et al., 2022) allowing us to study the spatially resolved properties of galaxies in these spectral lines. In comparison to the FoV of other time-expensive IFUs such as MaNGA (12′′- 32′′), SAMI (15′′), PPaK/CALIFA (60′′), and MUSE (60′′), the uniqueness of J-PLUS is its large FoV () offering a large contiguous observing area to understand the resolved structure of galaxies, galaxy groups, galaxy clusters, and thus the ability to probe their environment.
@@
The method of measuring H fluxes and star formation rates (SFRs) from J-PLUS photometry was first introduced by Vilella-Rojo et al. (2015) and Logroño-García et al. (2019). Vilella-Rojo et al. (2021) applied their method to estimate the star formation main sequence in the local Universe using J-PLUS DR1. Here, we aim to extend that work to validate the H flux extraction for spatially resolved studies using J-PLUS DR3. We selected galaxies with redshifts up to z 0.0165, ensuring that the J0660 NB filter covers the H emission line within this range. This leads to thousands of galaxies at D 70 Mpc within the 3192 deg2 observed area of J-PLUS DR3. At this redshift range, J-PLUS is able to resolve star-forming regions on the physical scale of several hundred parsecs ( 168 pc). We validated our method of building the emaps by comparing them with other well-known IFU observations of the same galaxies. This is the first of a series of works where we aim to use the H map to study the spatially resolved star formation main sequence of J-PLUS galaxies and ultraviolet (UV) and H studies of J-PLUS galaxies.
@@ 2.1 J-PLUS DR3
J-PLUS (Cenarro et al., 2014, 2019) is a multi band photometric survey designed to observe the northern sky in several thousand square degrees using the 80cm Javalambre Auxiliary Survey Telescope (JAST80) at OAJ, Teruel. It has 12 broad, intermediate, and narrow band optical filters with 2 deg2 FoV. The properties of J-PLUS filters are given in Table 1. After two data releases (DR1 on July 2018, and DR2 on July 2020), J-PLUS had its third data release in July 2022, containing 1642 J-PLUS fields covering 3192 deg2. The pixel scale of J-PLUS is 0.555′′/pixel. The point spread function (PSF) of J-PLUS observation is 1.1 arcsecs in the r band. The DR3 images that we have used here are available from the OAJ database 111https://archive.cefca.es/catalogues/jplus-dr3. The calibration and zeropoints are taken from López-Sanjuan et al. (2019, 2024).
@@ 2.2 PHANGS-MUSE
The MUSE (Bacon et al., 2010) is an IFS at the Very Large Telescope (VLT) of the European Southern Observatory (ESO) Paranal Observatory, Chile. It is composed of 24 identical integral-field units (IFUs). In its wide field mode (WFM; FoV of ) 90000 spectra are taken with a spatial resolution of 0.2′′/pixel covering a wavelength range of 4750 - 9350 Å . The spectral resolution of MUSE is 2.5 Å at 7000 Å and the spectral sampling of 1.25Å /slice.
@@ 2.3 CALIFA
CALIFA (Sánchez et al., 2012) is a large project to obtain spatially resolved spectra of 600 nearby galaxies with a size of 64′′ (diameter) in the redshift range of 0.005 z 0.03 using a PMAS/PPAK integral field spectrophotometer, mounted on the Calar Alto 3.5 m telescope. It utilises an hexagonal field-of-view (FoV) of . It has two observing modes: Intermediate spectral resolution (V1200, R1550) and low resolution (V500, R850). The CALIFA observations were carried out between 2010 and 2016.
We used version 2.3 from the extended data release (eDR) of CALIFA (Sánchez et al., 2023) that includes 895 galaxies. This version has significantly improved spatial resolution and image quality. The PSF is improved from 2.4 arcsec to 1 arcsec of full width at half-maximum (FWHM). The configuration includes the low-resolution setup (V500, R850), integration time of 900 s per pointing, with a wavelength range of 3745 and 7500 Å . The spatial sampling of the eDR CALIFA data is 0.5′′/pixels with a wavelength resolution of 2Å /slice. The eCALIFA cubes are downloaded from the CALIFA website 333https://ifs.astroscu.unam.mx. The downloaded CALIFA cubes are already corrected for Galactic extinction using the method by Schlegel et al. (1998) and Cardelli et al. (1989).
@@ 2.4 MaNGA
MaNGA (Bundy et al., 2015) is a program under the fourth phase of the Sloan Digital Sky Surveys (SDSS-IV; Blanton et al., 2017). MaNGA has mapped the kinematics and composition of 10,100 nearby galaxies in the redshift range of 0.01 z 0.15 with an average of z0.03 using 17 hexagonal shaped IFUs mounted on Sloan Foundation 2.5m telescope at Apache Point Observatory from 2014 to 2020. The IFU size ranges from 12 arcsecs (19 fibers) diameter to 32 arcsecs (127 fibers) with a wavelength coverage of 3600 - 10000 Å across 2700 deg2 and a spectral resolution of R2000. The medium spatial resolution of MaNGA is 2.54′′. The pixel scale of MaNGA is 0.5′′/pixel with a wavelength resolution of 0.834 Å /slice. The MaNGA cubes are downloaded from DR17 archive444https://data.sdss.org/sas/dr17/manga/spectro/redux/v3_1_1. The MaNGA fluxes are not corrected for Galactic extinction.
@@ 3 Sample selection
There are 176 galaxies cross-matched with J-PLUS DR3 and eCALIFA, with 78 galaxies in the redshift range of z0.0165. Among them, 20 galaxies are early-type (E/SO), nine galaxies show dust-lane features, and 48 galaxies show clear evidence of star formation activity (based on GALEX-UV and H + [NII] images) in the visual morphology and exhibit H emission line at 6562.8 Å rest frame wavelength in CALIFA spectra. Among the 48 galaxies, the [NII] line lies outside the filter window for redshifts greater than 0.014 for 10 galaxies. Above redshifts greater than 0.0165, the H line begins to fall outside the filter curve, and galaxies with such redshifts were excluded from the sample.
@@
The total sample includes 158 galaxies with 3D data cubes from either of the aforementioned IFUs. For the final analysis, we used 115 galaxies that exhibit H line in their spectra and have coverage of the full H line within the J0660 filter curve. Fig. 3 shows the coverage of J-PLUS DR3 with the distribution of the final sample selected from the cross-matched catalog. The information of the final cross-matched sample is given in Table A1.
@@ 4 The J-SHE Pipeline for building emaps of J-PLUS and IFUs
In this section, we describe the methodology followed to build the emaps and radial profile of galaxies, spectral energy distributions (SEDs), and spectra of emission line regions from J-PLUS and IFUs. The J-PLUS Spatially resolved H Emission line (J-SHE) pipeline is written in GNU Make mainly utilizing GNU Astronomy Utilities (Gnuastro; Akhlaghi & Ichikawa, 2015) and Python packages. Gnuastro is an official GNU package for the manipulation and analysis of astronomical data. Our pipeline consists of three modules (see Fig. 4). The first and second modules are dedicated to building emaps for J-PLUS (hereafter, J-IFU module) and IFUs (hereafter, IFU module) respectively. The third module generates the galaxy radial profiles, SEDs, and spectra of emission line regions from the emaps (hereafter, analysis module).